if steven can mix 20 drinks in 5 minutes.sue can mix 20 drinks in 10 minutes. jack can mix 20 drinks in 15 minutes.how much time will it take all three of them working together to mix the 20 drinks
Answer by ptaylor(2052)


Let x= amount of time it takes all three working together to mix 20 drinks
Steven mixes drinks at the rate of 20/5 or 4 drinks per min
Sue mixes drinks at the rate of 20/10 or 2 drinks per min
Jack mixes drinks at the rate of 20/15 or 1 1/3=4/3 drinks per min
Together, the three mixes drinks at the rate of 4+2+4/3 drinks per min or:
12/3 + 6/3 +4/3 = 22/3 drinks per min. So or equation to solve is:
(22/3)x=20 multiply each side by 3
22x=60 divide each side by 22
x=2.73 min------------------------------------ans
CK
Steven: 4*2.73=10.9 drinks
Sue: 2*2.73=5.46 drinks
Jack: (4/3)*2.73=3.64 drinks
10.9+5.46+3.63=20
20=20
Hope this helps--ptaylor